OpenStudy (anonymous):

find two numbers whose sum is 46 and whose product is a minimum

OpenStudy (anonymous):

You want two numbers x and y such that x + y = 46 and we want to minimize xy. Since x + y = 46, we have y = 46 - x, so \[ xy = x \cdot \left( 46 - x \right) = 46x - x^2. \] So we want the value of x so that this expression is minimized. Do you know how to do this?

OpenStudy (anonymous):

I understand how to do the equation, but I need a "minimized" answer, after you take the derivative you get the maximized answered of 23, how do you get the minimum ?

OpenStudy (anonymous):

23^2?

OpenStudy (anonymous):

is this the answer?

OpenStudy (anonymous):

its telling me 23 is wrong

OpenStudy (anonymous):

no 23^2=529

OpenStudy (anonymous):

that is also wrong

OpenStudy (dan815):

23^2 wud lead to the greatest product

OpenStudy (dan815):

something like 45.9999999999 * 0.0000000001

OpenStudy (anonymous):

thanks for trying everyone, im out

OpenStudy (dan815):

46 and 0 is

OpenStudy (anonymous):

also wrong

OpenStudy (dan815):

is there a restriction like positive non negative integer solutions?

OpenStudy (dan815):

no zero allowed

OpenStudy (dan815):

oh then lets make the product negative

OpenStudy (anonymous):

23 and 23 is a stationary point

OpenStudy (anonymous):

23 and 23 is the maximum, i need the minimum

OpenStudy (dan815):

yes yes

OpenStudy (dan815):

calm your buns

OpenStudy (swissgirl):

LMAO Reported ^

OpenStudy (dan815):

bahaha

OpenStudy (anonymous):

useless, most of you

OpenStudy (anonymous):

question is wrong. see given any number x, you can get another smaller x satisfying x+y=46

OpenStudy (dan815):

oh i see

OpenStudy (dan815):

-inf * inf+46

OpenStudy (dan815):

makes sense

OpenStudy (anonymous):

lox not inf, but tends to inf, this question was self made, not from any book

OpenStudy (dan815):

-inf * (inf+46)

OpenStudy (dan815):

just an argument to show why the product can be like -inf

OpenStudy (anonymous):

first take -100,100+46, then take -1000000000 , 10000000000+46, it decreases without any boundary

OpenStudy (anonymous):

just a waste of time, for an invalid question.

OpenStudy (dan815):

yes i just said this

OpenStudy (dan815):

i have a feeling the teacher meant to give bounds or something

OpenStudy (anonymous):

@vivzchic22 did you get what i meant to say? first take x=-100, y=100+46, then take x=-1000, y=1000+46, .... then take x=-1000000000, y=1000000000+46, the product is not bounded, it decreases wihtout any boundary, hence no minimum value

OpenStudy (dan815):